Fresh Co. purchased a new machine on July 2, 2019, at a total installed cost of $41,000. The machine has an estimated life of five years and an estimated salvage value of $6,600.

Required:

  1. Calculate the depreciation expense for each year of the asset's life using:
    1. Straight-line depreciation.
    2. Double-declining-balance depreciation.
  2. How much depreciation expense should be recorded by Fresh Co. for its fiscal year ended December 31, 2019, under each method? (Note: The machine will have been used for one-half of its first year of life.)
  3. Calculate the accumulated depreciation and net book value of the machine at December 31, 2020, under each method.
Answer #1

a) Depreciation expense 2019

Straight line Double decline
Depreciation expense (41000-6600/5)*6/12 = 3440 41000*40%*6/12 = 8200

b) Calculate following

Straight line Double decline
Accumulated depreciation Net book value Accumulated depreciation Net book value
2020 3440*3 = 10320 41000-10320 = 30680 41000*60%*40%*6/12+16400 = 21320 41000-21320 = 19680